OpenStudy (bohotness):

help have20 question lefdt

OpenStudy (bohotness):

Malia incorrectly graphed the equation y = x + 12 with a slope of . What should Malia's corrected graph look like? A. vertical B. more steep C. horizontal D. less steep

OpenStudy (bohotness):

@CausticSyndicalist @Hotchellerae21 @Teddyiswatshecallsme

OpenStudy (anonymous):

with slope, of what?

OpenStudy (anonymous):

where is the picture?

OpenStudy (bohotness):

s 3its 2/3x + 12 =with slope 3/2

OpenStudy (anonymous):

it should be more sttep as 2/3<1

OpenStudy (anonymous):

if she grathed with slope 2/3 and she need to grath with slope 1

OpenStudy (bohotness):

Malia incorrectly graphed the equation y = 2/3 + 12 with a slope of 3/2 What should Malia's corrected graph look like? A. vertical B. more steep C. horizontal D. less steep

OpenStudy (anonymous):

oh ok then le t me think

OpenStudy (anonymous):

marias grath should be less steep as 2/3<3/2

OpenStudy (bohotness):

ok

OpenStudy (bohotness):

Which of the following graphs shows a negative rate of change? .

OpenStudy (bohotness):

@ttb123456789

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

7b

OpenStudy (anonymous):

can you close this question and post new one to give me medal?